# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v10

@@ 6.2 Neighborhood Density
Decreasing Exteriors Lemma Suppose that $G$ is an oriented graph represented by a GLOVER, where there are no back arcs, with $u_{i}\in R_{i}$ and $v_{i+1}\in N^{+}(u_{i})\cap R_{i+1}$, where both $u_{i}$ and $v_{i+1}$ have the Decreasing Neighborhood Sequence Property (DNSP). Then, for all $z\in\text{ext}(u_{i},v_{i+1})$, we have $|\text{ext}(v_{i+1},z)|<|\text{ext}(u_{i},v_{i+1})|$. That is, the exterior degree of $v_{i+1}$ with respect to $u_{i}$ decreases for every second neighbor $z$ of $u_{i}$.
@@
Let us first restate our definitions. The DNSP say that a node’s second neighborhood is strictly less than its first neighborhood $|N^{++}(u)|<|N^{+}(u)|$
Recall that for an arc $(u_{i},v_{i+1})$, $\text{int}(u_{i},v_{i+1})=N^{+}(x)\cap N^{+}(v_{i+1})$.
Similarly, for that same arc $(u_{i},v_{i+1})$, $\text{ext}(u_{i},v_{i+1})=N^{++}(u_{i})\cap N^{+}(v_{i+1})$.
And for the arc $(v_{i+1},z)$, $\text{ext}(v_{i+1},z)=N^{++}(v_{i+1})\cap N^{+}(z$.
@@
Assume that we have a node $v_{i+1}\in R{i+1}$, whose interior degree doubles. We assumed there are no back arcs, which means that for all $\forall w\in\text{ext}(u_{i},v_{i+1})\cap\text{ext}(v_{i+1},z)=\emptyset$.
@@
For the sake of contradiction, suppose that there exists a $z\in\text{ext}(u_{i},v_{i+1})$ such that $|\text{ext}(v_{i+1},z)|\geq|\text{ext}(u_{i},v_{i+1})|$.
@@
We can break the degree of $v_{i+1}$ down into two components (remember there are no back arcs that would be a third component), its interior degree and its exterior degree $d^{+}(v_{i+1}=d|\text{int}(u_{i},v_{i+1})|+|\text{ext}(u_{i},v{i+1})|$. The Lemma 6.4 (Interior Degree Doubles) tells us that the node $v_{i+1}$ has its interior degree double. Thus, we have that if we can show that $v_{i+1}$’s exterior degree also doubles, it would lead to a contradiction because $v_{i+1}$ would have both its interior and exterior degree doubling, and thus its overall degree doubling, violating the DNSP.
@@
Our assumption for the sake of contradiction was that $z\in\text{ext}(u_{i},v_{i+1})$ such that $|\text{ext}(v_{i+1},z)|\geq|\text{ext}(u_{i},v_{i+1})|$. The node $v_{i+1}$’s out-degree will be impacted by $\text{ext}(u_{i},v_{i+1})$ and $\text{ext}(v_{i+1},z)$. This means that $z\in\text{ext}(u_{i},v_{i+1})=N^{++}(x)\cap N^{+}(v_{i+1})=R_{i+2}$
@@
If $|\text{ext}(v_{i+1},z)|\geq|\text{ext}(u_{i},v_{i+1})|$, then $|\text{ext}(v_{i+1},z)|+|\text{ext}(u_{i},v_{i+1})|\geq 2|\text{ext}(u_{i},v_{i+1})|$
We have two exterior sets on the left of the equation that we would like to sum, but we need to be sure that there is no node $w\in\text{ext}(v_{i+1},z)\cap\text{ext}(u_{i},v_{i+1})$.
@@
Consider the set $I=N^{++}(v_{i+1})\cap N^{+}(z)\cap N^{++}(u_{i})\cap N^{+}(v_{i+1})$. The set $I(x,v_{i+1},z)=\text{ext}(v_{i+1},z)\cap\text{ext}(u_{i},v_{i+1})$ is the intersection of those two sets.
Our job is to determine the cardinality of $I(u_{i},v_{i+1},z)$. Namely can there exist a $w\in I(u_{i},v_{i+1},z)$. $w\in I$ implies that there are two possibilities for $w$. Either $w\in N^{++}(u_{i})\cap N^{+}(v_{i+1})$ or $w\in N^{++}(v_{i+1})\cap N^{+}(z)$.
If $w\in N^{++}(u_{i})\cap N^{+}(v_{i+1})$, this would mean that there was an arc $(u_{i},w)$ across two rooted neighborhoods, skipping all nodes in the rooted neighborhood $R_{i+1}$. We do not allow that through our shortest path constructions though.
If $w\in N^{++}(v_{i+1})\cap N^{+}(z)$, then we have an arc from the rooted neighborhood $R_{i+2}$ where $z$ lies, to $R_{i+1}$ where $v_{i+1}$ and $w$ would be. This arc $(z,w)$ would represent a back arc though, which we assumed did not exist in this graph. This would mean that the exterior degree of $v_{i+1}$ is at least as large as the exterior degree of $u_{i}$.
@@
Thus, the assumption $|\text{ext}(v_{i+1},z)|\allowbreak\geq\allowbreak|\text{ext}(u_{i},v_{i+1})|$ leads to a contradiction, confirming that $|\text{ext}(v_{i+1},z)|\allowbreak<\allowbreak|\text{ext}(u_{i},v_{i+1})|$ for all $z\allowbreak\in\allowbreak\text{ext}(u_{i},v_{i+1})$.
Therefore, the lemma holds: for all $z\in\text{ext}(u_{i},v_{i+1})$, the exterior degree of $v_{i+1},z$ is strictly smaller than the exterior degree of $u_{i},v_{i+1}$. In other words, the size of the exterior neighbors decreases between consecutive neighborhoods.
This lemma is important because, in contrast to other areas in GLOVER representations, we actually see at a set-theoretic level how back arcs can disrupt the GLOVER data structure. It is not just dual paths, but it is in overlapping exterior sets. Once again, this must be addressed carefully. We cannot simply assume it away.
Section 7 (Back Arcs) does address this as well. Even though this is a different issue. it is still resolved by the no back arcs lemma we present there. When back arcs are present, we will not be able to present the oriented graph in a GLOVER. Instead, they will still lead directly to a degree doubling node. We will still be able to take advantage of splitting up a nodes degree equation. The notion of interior degree doubling has not been impacted by back arcs and we will be able to take advantage of that. These back arcs will still lead to second neighbors of the source node and a different path towards this doubling algorithm.
Example 6.3.
@@
The image above illustrates the situation that can cause the two exteriors on the left hand side in the proof in 6.5. The node $u_{i}$ has one out-neighbor, $v_{i+1}$. The node $v_{i+1}$ has two out-neighbors, $\{w,z\}$. This makes $\text{ext}(u_{i},v_{i+1})=(w,z)$. To calculate $\text{ext}(v_{i+1},z)$, we first calculate the nodes at distance two from $v_{i+1}$. We must traverse through the back arc, $(z,w)$. Then to traverse the node of distance one, we just traverse that same arc and see that $\text{ext}(v_{i+1},z)=\{w\}$.
Lemma 6.6.
|$|\{y\in\text{ext}(u,v)\mid u\in R_{i},\ v\in R_{i+1},\ y\in R_{i+2}\}|\leq d^{+}(v_{0})-i.$
Proof.
We proceed by induction on the distance $i$.
Base Case ($i=0$): The neighborhood $R_{0}=\{v_{0}\}$, and it’s out-neighbors form $R_{1}$, with $|R_{1}|=d^{+}(v_{0})$ by definition. Now consider $R_{2}$, which consists of nodes reachable by an arc from nodes in $R_{1}$. Let $v_{0}\in R_{0}$, $v\in N^{+}(v_{0})\subseteq R_{1}$, and define $\text{ext}(v_{0},v):=N^{++}(v_{0})\cap N^{+}(v)$. From DNSP,
$|N^{++}(v_{0})|<|N^{+}(v_{0})|\leq d^{+}(v_{0}),$
so
$|\text{ext}(v_{0},v)|\leq d^{+}(v_{0})-1.$
Therefore,
|$|\{y\in\text{ext}(v_{0},v)\mid y\in R_{2}\}|\leq d^{+}(v_{0})-1.$
Inductive Hypothesis: Assume that for some $i\geq 0$,
|$|\{y\in\text{ext}(u,v)\mid u\in R_{i},\ v\in R_{i+1},\ y\in R_{i+2}\}|\leq d^{+}(v_{0})-i.$
Inductive Step: Consider $u\in R_{i+1}$, $v\in N^{+}(u)\subseteq R_{i+2}$, and we will examine $\text{ext}(u,v):=N^{++}(u)\cap N^{+}(v)$.
From Lemma 6.5 (Decreasing Exteriors Lemma), we know:
$|\text{ext}(u,v)|<|\text{ext}(u^{\prime},u)|,$
for any parent $u^{\prime}\in R_{i}$, $u\in R_{i+1}$.
By the inductive hypothesis,
$|\text{ext}(u^{\prime},u)|\leq d^{+}(v_{0})-i.$
Therefore,
$|\text{ext}(u,v)|<d^{+}(v_{0})-i\Rightarrow|\text{ext}(u,v)|\leq d^{+}(v_{0})-(i+1).$
What this has established through the DNSP is that the rooted neighborhoods are decreasing in size as we move further away from the minimum out-degree node $v_{0}$. We spoke of the DNSP being a bottleneck, and here is another example where, across the entire graph, we are actually seeing it restrict the sizes of rooted neighborhoods. DNSP is a local bottleneck that induces a global contraction of the neighborhood layers in the GLOVER.
We began with the assumption that the Decreasing Neighborhood Sequence Property (DNSP) established that for a node $u$, the size of its first out-neighborhood is strictly larger than its second out-neighborhood $|N^{+}(u)|>|N^{++}(u)|$. We have now witnessed the emergence of this principle at the level of entire rooted neighborhoods within the GLOVER structure. Lemma 6.6 (DNSP Impact on Rooted Neighborhood Size) shows that consecutive neighborhoods must strictly decrease in size $(|R_{i+1}|<|R_{i}|$ for $i\geq 1)$. This demonstrates DNSP rising from a local node-level constraint to a global feature of the graph’s partitioning. This highlights the penetrating and fundamental nature of the DNSP. It is felt at every node of the graph, impacting not only the individual nodes but the hierarchies above them as well. Because the GLOVER is organized in this way, it merges the power of the DNSP to truly organize these levels. This rooted neighborhood-level constraint of the DNSP sets the stage for the next level of understanding between the implications for the existence of back arcs. These back arcs hold weight in whether GLOVER is just another data structure or if it can help solve the Seymour Second Neighborhood Conjecture.
We have shown $|R_{i+1}|<|R_{i}|$, which means $|R_{i+1}|\leq|R_{i}|-1$. Now that we know that the rooted neighborhoods decrease in size, we can quantify the bounds on these sizes based on the distances from the minimum out-degree node. This will help with things like back arcs and the ability to generalize the above lemma.
Lemma 6.7.
$|R_{0}|=1,|R_{1}|\leq\delta,|R_{i}|\leq\delta-(i-1)$
Proof.
We proceed by induction on the distance $i$ of rooted neighborhoods $R_{i}$.
Base cases:
  • $R_{0}=\{v_{0}\}$, so $|R_{0}|=1$ by definition.
  • $R_{1}=N^{+}(v_{0})$, so $|R_{1}|=\delta$, again by definition.
The DNSP says $|N^{+}(v_{0})|>|N^{++}(v_{0})|$. All of the out-neighbors of $v_{0}$ are in the rooted neighborhood $R_{1}$ and every second neighbor of $v_{0}$ will be in the rooted neighborhood $R_{2}$. By Lemma 6.2 (Interior Load Balancing), these children $u\in R_{1}$ help prevent $v_{0}$ from becoming a degree doubling node by forming a cycle in $R_{1}$. This bounds the number of external neighbors in $R_{2}$ and thus the size of $|R_{2}|$.
Inductive hypothesis: Assume that for some $k<i$, we have:
$|R_{k}|\leq\delta-(k-1),\quad\text{and for all }u\in R_{k}:d^{+}_{\text{ext}}(u)<\delta-(k-1),\quad d^{+}_{\text{int}}(u)>k-1.$
Inductive step: We apply Lemma 6.6 (DNSP Impact on Neighborhood Size), which implies:
$|R_{i}|<|R_{i-1}|\leq\delta-(i-2)\Rightarrow|R_{i}|\leq\delta-(i-1).$
Now we can take any $u\in R_{i}$. Since back arcs are excluded and $u\in R_{i}$, $N^{+}(u)\subseteq R_{i}\cup R_{i+1}$. Thus we can calculate exterior neighbors by:
$d^{+}_{\text{ext}}(u)\leq|R_{i+1}|$. But from above, $|R_{i+1}|\leq\delta-i$, so:
$d^{+}_{\text{ext}}(u)<\delta-(i-1).$
Likewise, DNSP implies that the number of interior neighbors can be calculated by :
$d^{+}(u)>|N^{++}(u)|\Rightarrow d^{+}_{\text{int}}(u)>d^{+}_{\text{ext}}(u),$
So we can use the bound on exterior degree:
$d^{+}_{\text{int}}(u)>\delta-(i-1)-1=\delta-i.$
Since $\delta$ is fixed, and $i\geq 1$, we conclude:
$d^{+}_{\text{int}}(u)>i-1.$
We conclude that the inductive structure holds for all $i$, giving:
$|R_{i}|\leq\delta-(i-1),$
and for each $u\in R_{i}$:
$d^{+}_{\text{ext}}(u)<\delta-(i-1),\quad d^{+}_{\text{int}}(u)>i-1.$
What we saw from 6.6 (DNSP Impact on Rooted Neighborhood Size) is that the rooted neighborhoods decreased in size. This decrease must be at least by one at each neighborhood. What 6.7 (Formula for Rooted Neighborhood Size) does is combine this into a worse case calculation size for each rooted neighborhood. In addition, this allows us to calculate the worst case sizes for interior and exterior neighborhood sizes. This gives us a way to evaluate a node’s second neighborhood against these rooted neighborhood sizes, as a comparison for degree-doubling nodes.
@@
Let us first restate our definitions. The DNSP says that a node’s second neighborhood is strictly less than its first neighborhood $|N^{++}(u)|<|N^{+}(u)|$. For an arc $u\rightarrow v$ with $u\in R_{i-1}$ and $v\in R_{i}$, define: The interior out-degree of $v$ with respect to $u$ as $|\text{int}(u,v)|:=|N^{+}(u)\cap N^{+}(v)|$ The exterior out-degree of $v$ with respect to $u$ as $|\text{ext}(u,v)|:=|N^{++}(u)\cap N^{+}(v)|$
@@
By Lemma 6.4 (Interior Degrees Doubling), interior degrees grow at least linearly along the path from the minimum degree node to $u_{i+1}$. If exterior degrees also decrease adequately, total out-degree would grow beyond the DNSP bounds.
Lemma 6.5 (Decreasing Exteriors) states that along the path from $v_{0}$ to $u_{i}$ (the parent node of $u_{i+1}$, exterior degrees must strictly decrease. Formally, for nodes $u_{k}\in R_{k}$ and $u_{k+1}\in R_{k+1}$, we have the following equation:
@@
As mentioned above, our initial lemma 6.2 exists in literature on the Seymour Conjecture [1]. The truly critical insight lies in the generalization of the load-balancing principle we have uncovered within the GLOVER data structure.
@@
Every node, at every level of the GLOVER, must participate in effectively distributing outgoing connections to the subsequent rooted neighborhood. This collective load balancing is essential to ensure that predecessors at earlier levels do not experience a doubling of their degree. The failure of any node within any rooted neighborhood to adequately contribute to this distribution inevitably leads to degree doubling of a predecessor.
This generalized principle underscores the intricate and interconnected nature of the GLOVER, highlighting that the avoidance of degree doubling is a global constraint dependent on the cooperative behavior of nodes across all rooted neighborhoods, far beyond the easily observable base cases.
Example 6.4.
0
$\Rightarrow$ 111122223333444455556666777788889999101010101111111112121212131313131414141415151515161616161717171718181818191919192020202021212121222222222323232324242424H
[…]
Figure 12: A graph visualization illustrating the Decreasing Neighborhood Sequence Property with a minimum out-degree of 9. Each radial layer $R_{i}$ represents the set of nodes at increasing distance from the root $v_{0}$. Nodes are shaded progressively darker to reflect increasing interior out-degree within their respective rooted neighborhoods. Arrows indicate interior arcs between nodes. Note that neither node 0 nor nodes in the two outer layers experience degree doubling.
Example 6.4 shows some of the patterns that emerge with these interior degrees in the SSNC. Starting from the outside in, going inward represents the order of rooted neighborhoods as they increase in distance from $v_{0}$. This example has visualized only the interior arcs, as the exterior arcs would complicate this visualization. However, just as in previous examples, all nodes in a previous rooted neighborhood are mapped to all nodes in the next rooted neighborhood.
Example 6.5.
@@
What we also saw in this section, though, is that as this load is passed from parent to child the size of these neighborhoods is reduced. This leads to a sequence of rooted neighborhoods that are getting smaller and smaller, but these smaller neighborhoods are expected to hold more nodes inside them. Obviously there is a collision that is inevitable. We need to ensure that other things like back arcs do not get in the way. The Decreasing Neighborhood Sequence Property (DNSP) says that we should continuously have these rooted neighborhoods decrease forever. Any time that they stop decreasing, we will arrive at a node, or set of nodes, whose degree doubles. This would be a violation of the DNSP assumption and a contradiction.
6.3 Representation
The more popular representations of graphs are adjacency matrices, array-lists, and edge lists. These suffer from the same limitations of graphs as mentioned above; they scatter the nodes along the two-dimensional plane without attempts to group them into an ordering. The Graph Level Order (GLOVER) improves this by partitioning the nodes in a reasonable way. For the SSNC, that reasonable way was the distance metric. To represent graphs in a GLOVER, we can use any data serialization that can encode graphs, including JSON, YAML, and XML. There needs to be a new level of data: rooted neighborhoods, to go along with the standard data serialization. The neighborhoods will be given an ID associated with their distance from the minimum out-degree node. The nodes will then be assigned to their unique neighborhoods. Arcs will also be determined to be interior, exterior, or back, depending on the endpoints of the nodes of the arcs.
Node Targets Neighborhood
0 1, 2, 3 $R_{0}$
1 2, 4, 5, 6 $R_{1}$
2 3, 4, 5, 6 $R_{1}$
3 1, 4, 5, 6 $R_{1}$
4 5, 7, 8 $R_{2}$
5 6, 7, 8 $R_{2}$
6 7, 8, 1 $R_{3}$
7 8, 6, 1 $R_{3}$
8 6, 7, 1 $R_{3}$
Table 1: This figure illustrates a JSON representation of a GLOVER data structure in array-list representations. Each node has a list of targets, along with its assigned neighborhood. Not shown are the nodes in neighborhood $R_{4}$. In this JSON format, each key represents a node ID, and its value is an object containing ’targets’ (a list of its out-neighbors) and ’neighborhood’ (the rooted neighborhood it belongs to, $R_{0}$ being the minimum out-degree node). This structure explicitly groups nodes by their distance from $v_{0}$.
Example 6.6.
@@ 6.3 Representation
Table 2: A JSON representation of a GLOVER data structure with an irregular $R_{1}$ interior degree but exhibiting load balancing. The interior degrees in $R_{1}$ vary (irregularity), yet sufficient internal connections within $R_{1}$ maintain overall balance, preventing degree doubling in $R_{0}$.

### v12

@@ 6.2 Neighborhood Density
Exterior Degree Doubling Let $G$ be an oriented graph represented by a Graph Level Order with root node $v_{0}$. Further , assume $G$ has no back arcs. Suppose $u_{i}\in R_{i}$ has its interior out-degree double and $\text{ext}(u_{i-1},u_{i})=R_{i+1}$ Then if $|R_{i+1}|\geq|R_{i+2}|$, $u_{i}$ will have its exterior out-degree double as well, causing the overall out-degree of $u_{i}$ to double.
@@
Let us begin by restating our definitions. For a node $u_{i}$, we define the interior out-degree with respect to a parent $u_{i-1}$ as
$|\text{int}(u_{i-1},u_{i})|.$
Similarly, we define the exterior out-degree with respect to that same parent as
$|\text{ext}(u_{i-1},u_{i})|.$
By the assumption that $u_{i}$’s interior out-degree doubles, we are assuming that the number of second interior neighbors of $u_{i}$ with respect to $u_{i-1}$ is at least as large as its number of first interior neighbors. That is,
$|N^{++}_{\text{int}}(u_{i})|\geq|N^{+}_{\text{int}}(u_{i})|.$
By Lemma 5.1 (Partition of Node’s Degree), and the lack of back arcs in $G$, the behavior of a node’s degree doubling is dependent on its interior out-degree doubling and its exterior out-degree doubling. Because these sets are disjoint, they act independently of one another. Thus, if we can also show that $u_{i}$’s exterior out-degree doubles, it will show that $u_{i}$’s overall degree doubles.
By assumption, the nodes of $\text{ext}(u_{i-1},u_{i})$ are in $R_{i+1}$. These are the first exterior out-neighbors of $u_{i}$. The nodes in $R_{i+2}$ are the first out-neighbors of the nodes in $R_{i+1}$, thus they are the second exterior out-neighbors of $u_{i}$. The assumption that $|R_{i+1}|=|R_{i+2}|$ says that the node $u_{i}$ has an equal number of first exterior out-neighbors and second exterior out-neighbors. That is,
$|N^{++}_{\text{ext}}(u_{i})|=|N^{+}_{\text{ext}}(u_{i})|.$
This makes $u_{i}$ satisfy the degree doubling condition for its exterior degree. We have already shown that $u_{i}$ satisfies the degree doubling condition for its interior degree, so $u_{i}$ satisfies the degree doubling condition for its overall degree.
This completes the proof. ∎
Lemma 6.5 is the counterpart of Lemma 6.4. It will help us with our traversal algorithm, as we will walk through the graph and seek conditions that will make a node’s exterior degree double. This lemma formalizes that when we have two neighborhoods that are equal in size, there is a node whose degree doubles.
Lemma 6.6.
Decreasing Exteriors Lemma Suppose that $G$ is an oriented graph represented by a Graph Level Order, where there are no back arcs, with $u_{i}\in R_{i}$ and
$v_{i+1}\in N^{+}(u_{i})\cap R_{i+1},$
where both $u_{i}$ and $v_{i+1}$ have the Decreasing Neighborhood Sequence Property (DNSP) and the node $v_{i+1}$ has its interior degree double. Then, for all
$z\in\text{ext}(u_{i},v_{i+1}),$
we have
$|\text{ext}(v_{i+1},z)|<|\text{ext}(u_{i},v_{i+1})|.$
That is, the exterior degree of $v_{i+1}$ with respect to $u_{i}$ decreases for every second neighbor $z$ of $u_{i}$.
Proof.
@@
Let us first restate our definitions. The DNSP say that a node’s second neighborhood is strictly less than its first neighborhood
$|N^{++}(u)|<|N^{+}(u)|.$
Recall that for an arc $u_{i}\to v_{i+1}$,
$\text{int}(u_{i},v_{i+1})=N^{+}(x)\cap N^{+}(v_{i+1}).$
Similarly, for that same arc $u_{i}\to v_{i+1}$,
$\text{ext}(u_{i},v_{i+1})=N^{++}(u_{i})\cap N^{+}(v_{i+1}).$
And for the arc $v_{i+1}\to z$,
$\text{ext}(v_{i+1},z)=N^{++}(v_{i+1})\cap N^{+}(z).$
@@
Assume that we have a node $v_{i+1}\in R_{i+1}$, whose interior degree doubles. We assumed there are no back arcs, which means that for all
$\forall w\in\text{ext}(u_{i},v_{i+1})\cap\text{ext}(v_{i+1},z)=\emptyset.$
@@
For the sake of contradiction, suppose that there exists a $z\in\text{ext}(u_{i},v_{i+1})$ such that
$|\text{ext}(v_{i+1},z)|\geq|\text{ext}(u_{i},v_{i+1})|.$
@@
We can break the degree of $v_{i+1}$ down into two components (remember there are no back arcs that would be a third component), its interior degree and its exterior degree
$d^{+}(v_{i+1})=|\text{int}(u_{i},v_{i+1})|+|\text{ext}(u_{i},v_{i+1})|.$
By assumption, the node $v_{i+1}$ has its interior degree double. Thus, we have that if we can show that $v_{i+1}$’s exterior degree also doubles, it would lead to a contradiction because $v_{i+1}$ would have both its interior and exterior degree doubling, and thus its overall degree doubling, violating the DNSP.
Our assumption for the sake of contradiction was that $z\in\text{ext}(u_{i},v_{i+1})$ such that
$|\text{ext}(v_{i+1},z)|\geq|\text{ext}(u_{i},v_{i+1})|.$
@@
The node $v_{i+1}$’s out-degree will be impacted by $\text{ext}(u_{i},v_{i+1})$ and $\text{ext}(v_{i+1},z)$. This means that
$z\in\text{ext}(u_{i},v_{i+1})=N^{++}(x)\cap N^{+}(v_{i+1})=R_{i+2}$
@@
If
$|\text{ext}(v_{i+1},z)|\geq|\text{ext}(u_{i},v_{i+1})|,$
then
$|\text{ext}(v_{i+1},z)|+|\text{ext}(u_{i},v_{i+1})|\geq 2\cdot|\text{ext}(u_{i},v_{i+1})|$
We have two exterior sets on the left of the equation that we would like to sum, but we need to be sure that there is no node
$w\in\text{ext}(v_{i+1},z)\cap\text{ext}(u_{i},v_{i+1}).$
@@
Consider the set
$I=N^{++}(v_{i+1})\cap N^{+}(z)\cap N^{++}(u_{i})\cap N^{+}(v_{i+1}).$
The set
$I(x,v_{i+1},z)=\text{ext}(v_{i+1},z)\cap\text{ext}(u_{i},v_{i+1})$
is the intersection of those two sets.
Our job is to determine the cardinality of
$I(u_{i},v_{i+1},z).$
Namely can there exist a
$w\in I(u_{i},v_{i+1},z).$
$w\in I$ implies that there are two possibilities for $w$. Either
$w\in N^{++}(u_{i})\cap N^{+}(v_{i+1})$
or
$w\in N^{++}(v_{i+1})\cap N^{+}(z).$
If
$w\in N^{++}(u_{i})\cap N^{+}(v_{i+1}),$
This would mean that there was an arc $u_{i}\to w$ across two rooted neighborhoods, skipping all nodes in the rooted neighborhood $R_{i+1}$. We do not allow that through our shortest path constructions though.
If
$w\in N^{++}(v_{i+1})\cap N^{+}(z),$
then we have an arc from the rooted neighborhood $R_{i+2}$ where $z$ lies, to $R_{i+1}$ where $v_{i+1}$ and $w$ would be. This arc $z\to w$ would represent a back arc though, which we assumed did not exist in this graph. This would mean that the exterior degree of $v_{i+1}$ is at least as large as the exterior degree of $u_{i}$.
@@
Thus, the assumption
$|\text{ext}(v_{i+1},z)|\allowbreak\geq\allowbreak|\text{ext}(u_{i},v_{i+1})|$
leads to a contradiction, confirming that
$|\text{ext}(v_{i+1},z)|\allowbreak<\allowbreak|\text{ext}(u_{i},v_{i+1})|$
for all
$z\allowbreak\in\allowbreak\text{ext}(u_{i},v_{i+1}).$
Therefore, the lemma holds: for all $z\in\text{ext}(u_{i},v_{i+1})$, the exterior degree of $v_{i+1},z$ is strictly smaller than the exterior degree of $u_{i},v_{i+1}$. In other words, the size of the exterior neighbors decreases between consecutive neighborhoods. ∎
This lemma is important because, in contrast to other areas in Graph Level Order representations, we actually see at a set-theoretic level how back arcs can disrupt the Graph Level Order. It is not just dual paths, but it is in overlapping exterior sets. Once again, this must be addressed carefully.
Section 7 (Back Arcs) does address this as well. Even though this is a different issue. It is still resolved by the no back arcs lemma we present there. When back arcs are present, we will not be able to present the oriented graph in a Graph Level Order. Instead, they will still lead directly to a degree doubling node. We will still be able to take advantage of splitting up a nodes degree equation. The notion of interior degree doubling has not been impacted by back arcs and we will be able to take advantage of that. These back arcs will lead to second neighbors of the source node and a different path towards this doubling algorithm.
@@
Figure 11: Illustration of a back arc causing exterior neighbors to have overlap. The arc $u_{i}\to v_{i+1}$ has exterior neighbors $w\text{ and }z$. The arc $v_{i+1}\to z$ has one exterior neighbor, $w$. We see that the node $w$ is in both sets.
@@
Figure 11 illustrates the situation that can cause the two exteriors on the left hand side in the proof in 6.6. The node $u_{i}$ has one out-neighbor, $v_{i+1}$. The node $v_{i+1}$ has two out-neighbors, $\{w,z\}$. This makes $\text{ext}(u_{i},v_{i+1})=\{w,z\}$. To calculate $\text{ext}(v_{i+1},z)$, we first calculate the nodes at distance two from $v_{i+1}$. Traversing through the back arc, $z\to w$. Then to traverse the node of distance one, we just traverse that same arc and see that $\text{ext}(v_{i+1},z)=\{w\}$.
More importantly though, what Lemma 6.6 (Decreasing Exteriors) showcases is that as we proceed on a path through exterior nodes from a minimum degree node $v_{0}$, the sizes of these exterior sets must decrease, as long as there are no back arcs along the path. We were first able to partition a node’s degree into interior and exterior degree by Lemma 5.1 (Partition of Node’s Degree). The point of Lemma 6.4 (Interior Degree Double) was to show that once this is done, the interior portion of this degree will always double. Now, this lemma and the DNSP are the foundations of our traversal algorithm through exterior neighbors.
For a more concrete example, consider the node $v_{0}$. This lemma is saying that the size of each exterior set we traverse, first through a node in $R_{1}$, then a node in $R_{2}$, etc, must be bounded first by $\delta-1$, then by $\delta-2$, continuing on throughout the graph. As we see, this is a much stronger lemma for graph traversal than Lemma 6.3 (Exterior Load Balancing).
@@
Let us first restate our definitions. The DNSP says that a node’s second neighborhood is strictly less than its first neighborhood $|N^{++}(u)|<|N^{+}(u)|$. For an arc $u\to v$ with $u\in R_{i-1}$ and $v\in R_{i}$, define: The interior out-degree of $v$ with respect to $u$ as
$|\text{int}(u,v)|:=|N^{+}(u)\cap N^{+}(v)|.$
The exterior out-degree of $v$ with respect to $u$ as
$|\text{ext}(u,v)|:=|N^{++}(u)\cap N^{+}(v)|.$
@@
Lemma 6.6 (Decreasing Exteriors) states that because we have interior degree at least $k$ for every node $u_{k}\in R_{k}$ along the path from $v_{0}$ to $u_{i}$ (the parent node of $u_{i+1}$), exterior degrees must strictly decrease as we traverse outwards. Formally, for nodes $u_{k}\in R_{k}$ and $u_{k+1}\in R_{k+1}$, we have the following equation:
@@
As mentioned above, our initial lemma 6.2 exists in literature on the Seymour Conjecture [6]. The truly critical insight lies in the generalization of the load-balancing principle we have uncovered within the Graph Level Order.
@@
Every node, at every level of the Graph Level Order, must participate in effectively distributing outgoing connections to the subsequent rooted neighborhood. This collective load balancing is essential to ensure that predecessors at earlier levels do not experience a doubling of their degree. The failure of any node within any rooted neighborhood to adequately contribute to this distribution inevitably leads to degree doubling of a predecessor.
This generalized principle underscores the intricate and interconnected nature of the Graph Level Order, highlighting that the avoidance of degree doubling is a global constraint dependent on the cooperative behavior of nodes across all rooted neighborhoods, far beyond the easily observable base cases.
Moreover, the inductive step made use of Lemma 6.4 (Interior Degree Doubling) and Lemma 6.6 (Decreasing Exteriors) to reach a contradiction. Now that we have an assumption of each node needing to support, not only their parents but their ancestors as well through higher degrees in further rooted neighborhoods, we can see the usage of the constructive 6.4 mapping algorithm. We know that there will always exist feasible cycle mappings for these nodes, no matter how large the rooted neighborhood. And because we know cycles immediately lead to degree-doubling nodes, this solves the interior degree case in $O(1)$ complexity. What remains is the traversal algorithm through the exterior nodes and back arcs.
@@
Table 3: A JSON representation (in table form) of a Graph Level Order with an irregular $R_{1}$ interior degree but exhibiting load balancing. The interior degrees in $R_{1}$ vary (irregularity), yet sufficient internal connections within $R_{1}$ maintain overall balance, preventing degree doubling in $R_{0}$.